# Superuniversal Statistics with Topological Origins for non-Hermitian Scattering Singularities

### v1

@@ III.1 Fundamental Singularities - Statistics
In panels (a-b) of Fig. 3, the color gradient corresponds to the absorption strength $\eta$, and there is a systematic, though small, vertical offset of the PDFs based on the ensemble value of $\eta$. This leads to different abundances of extreme values of $\mathcal{S}$. It is clear that the degree of loss is the strongest predictor of singularity density in parameter space. For any given quantity, the PDFs from ensembles with higher loss turn over to the power law tail at smaller values of the quantity and therefore have a lower overall probability at the extreme values associated with the scattering singularities. This agrees with the finding in [53], that the time delay PDFs from ensembles with larger $\eta$ had the power law tails onset earlier, resulting in smaller probabilities at extreme values, hence fewer singularities in systems with more uniform absorption.
@@ III.2 Singularity Dependence on Absorption
Figure 4: Singularity density as a function of loss in (a) experimental ensembles and (b) RMT simulation. Overall trend of reduced singularity density with increasing loss is easier to see in (b) as the experimental systems have more differences than just $\eta$ which cannot all be normalized out.
It might seem counterintuitive that singularity abundance would decrease with increasing absorption since singularities are zeros of the complex wave fields and absorption should suppress the fluctuation magnitude of the waves. However, singularities, as stated before, are primarily the result of interference effects of superpositions of random waves. In the presence of large amounts of dissipation, highly attenuated waves will not experience much destructive interference because the modes become more homogeneous and less random.
In Fig. 4(a) we show the singularity mode density, defined as the number of singularities per mode in the ensemble, of the fundamental singularities EP-2, CPA, RSM-1, and TSM-1s (or equivalently TSM-2s for reciprocal $\beta=1$ systems) found in our experimental ensembles as a function of $\eta$. As discussed in Sec. II, the mean mode spacing $\Delta$ is system size and frequency dependent, so $\eta$ is correlated to $\mathcal{D}$ in our experimental systems. Also, the techniques used to create ensembles are different for each $\mathcal{D}$ so despite the great effort made to create diverse ensembles, there is some systematic dependece for how ’ergodically’ we can explore the phase space of a given scattering environment based on what tunable parameters we have available. However, we still see a trend of lower singularity density with increasing $\eta$, agreeing with the complex time delay PDFs in Refs. [52, 53] that show tail behavior at smaller time delay values for systems with larger $\eta$. We performed RMT simulations using a Gaussian Orthogonal Ensemble (GOE) of 600 random matrices of size $10^{5}\times 10^{5}$ to generate scattering matrices following the procedure detailed in Appendix C, Ref. [142] and Appendix A of Ref. [141]. We use the same set of Hamiltonians and vary only the loss $\eta$ between ensembles, resulting in Fig. 4(b).
@@
Figure 5: (a) Locations of scattering zeros in the ($\omega,\nu$) parameter space of an RMT simulation with $\eta=0$. (b) Locations of scattering zeros in the ($\omega,\nu$) parameter space of an RMT simulation using the same effective Hamiltonian and mode-channel coupling as in (a) but with $\eta=0.006$. Dashed red and black lines show the $\eta=0$ locations of TSM-2 and $\delta R_{12}=0$ which have become points with the introduction of finite absorption. (c-d) Phase of $S_{+i}$ and $S_{-i}$ from simulation in (b). Black circles surround EPs (phase singularities) and arrows show direction of phase winding.
Note that at zero loss (i.e. a Hermitian system) certain singularities are forbidden, namely CPAs and EPs. Scattering parameter zeros (RSMs and TSMs) still occur because the elements of the $S$-matrix remain subunitary, as do symmetric reflection points and reciprocal points in non-reciprocal systems. However, the singularities can have different characteristics in a Hermitian system. Because we cannot experimentally measure a lossless system, we turn once more to RMT simulation. We consider the case of a reciprocal system with two channels ($\beta=1,M=2$) for simplicity using the model described in Ref. [91] Appendix H. This model allows for Hamiltonian $H_{0}$ with a tunable parameter $\nu$ with the equation
@@
where $H_{1},H_{2}$, and $H_{3}$ are all standard GOE random matrices. A scattering matrix that depends on both the parameter $\nu$ and a frequency $\omega$ can be calculated using coupled mode theory [147]:
@@
In Fig. 5(b) we show the resulting singularities when we recalculate the $S$-matrix using the same model and parameters as in (a), but now with a small uniform attenuation $\eta=0.006$. We find that when a DP is perturbed in a way that introduces non-Hermiticity, it transforms into two EP-2s. This is true for infinitesimally small amounts of loss, the smallest attempted in the simulations was $\eta=6\times 10^{-9}$, which can be seen in Appendix D. The charge of an EP-2 in a $M=2$ system is determined by whether the EP is a complex zero of $S_{+i}:=S_{11}-S_{22}-2i\sqrt{S_{12}S_{21}}$ (+i EPs marked in dark blue) or $S_{-i}:=S_{11}-S_{22}+2i\sqrt{S_{12}S_{21}}$ (-i EPs marked in light blue) [91], the phase of which are shown in panels (c) and (d), respectively. The black circles and winding arrows show that each DP splits into a pair of EP-2s with opposite winding and charge, such that the overall value of both these quantities is conserved when uniform loss is introduced.
In contrast, CPAs require some finite amount of loss to bring an $S$-matrix eigenvalue from the complex unit circle to zero. Exactly how much loss is required is a nuanced question since the trajectory of scattering eigenvalues in the complex plane is not linear. A lower bound, if it exists, should in principle be calculable, but is unknown to us. In the region of parameter space for the Hamiltonian used in Fig. 5, at $\eta=0.0006$ no CPAs could be found, however at $\eta=0.006$ a pair with opposite winding had already been created, marked by the green points in panel (b). With the addition of loss, not only have two CPAs appeared and the two DPs split into four EPs, but also the RSM-2s separated into RSM-1s (the pair corresponding to $|S_{22}|=0$ have already annihilated by $\eta=0.006$ so only two remain), and the $\delta R_{12}$ and TSM-2 lines have broken up into points.
@@ III.3 Higher Order Singularities - Statistics
Finally, we consider the question of higher order singularities which are composites of multiple fundamental singularities. Through the use of tunable perturbers, it is possible to make independent singularities coincident in parameter space, which results in novel applications such as the 50:50 IQ power splitter described in [91] using a coincident CPA-EP-2. The intersection of complex zeros of different fields has no topological stability, and the singularities can move apart under arbitrarily small perturbation. Consequently, we do not find a universal power law for all PDFs of functions that diverges at a singularity of order $P>2$, for any given $P$.
Fig. 6(a) demonstrates the experimental PDFs of $\frac{\langle|\mathcal{S}|\rangle}{|\mathcal{S}|}$ for $\mathcal{S}=1-|C|+|\text{det}S|$ which diverge at a combined CPA-EP-2 of the $S$-matrix. The different curves come from 6 ensembles with different values for the parameters $\mathcal{D},M,\beta,\eta$. The bright pink curve, which comes from the lossiest system measured ($\eta=145$), has a power law fit of $-3.49$ as demonstrated by the dashed blue line, while the bright yellow curve which comes from the system with the least absorption ($\eta=1.8$) has a power law fit of $-4.27$, demonstrated by the dashed black line. The four other curves at intermediate values of uniform attenuation fill out the space in between with slopes bounded by those two values. This shows that even though $\mathcal{P}(\frac{\langle|\text{det}S|\rangle}{|\text{det}S|})$ and $\mathcal{P}(\frac{\langle 1-|C|\rangle}{1-|C|})$ have universal power laws, the power law of a PDF of a higher order combined singularity is system detail-dependent. It is interesting to note that the steepest power law comes from the ensemble with the lowest loss and increasing absorption makes the PDF drop slower. This implies that although the density of individual singularities decreases with loss, their intersections become more common. It is currently unclear why this would be the case.
@@
A further statement should be made at this stage about higher order singularities of the scattering matrix. The fundamental singularities being combined are of not wholly independent complex scalars. CPAs are zeros of $\text{det}S$, but EPs also depend on the principal invariants of the $S$-matrix, one of which is its determinant. Similarly, $S_{xx}$ and $S_{xy}$ have inherent correlations which can depend on many factors, including the number of channels $M$ or loss $\eta$, as well as more complicated issues not considered in this paper such as imperfect mode-channel coupling. It is possible that truly universal statistics exist for composites of fully independent singularities, but there are not such higher order singularities for a physical scattering matrix.
@@ IV Conclusion
We have examined the statistical properties of $S$-matrix singularities by generalizing the concept of singularity speckle patterns to arbitrary two dimensional parameter spaces and any complex scalar function that describes wave phenomena involving complicated scattering. The statistics discussed come from experimental microwave resonant systems with the number of channels ranging from $M=1$ to $M=4$. Both systems with and without Lorentz reciprocity were considered, as well as a wide range of uniform absorption ($1.8\leq\eta\leq 145$). Supporting RMT simulations were conducted with results that agree with the empirical data, showing these findings are generic to all wave scattering systems, such as acoustic, optical, and photonic resonators. Fundamental singularities, which are topologically-protected vortices of complex scalar fields in two-dimensional parameter space, have a superuniversal statistical rule that a function which diverges at a singularity has a PDF with a -3 power law tail. These PDFs can then be used to estimate singularity abundance based on system parameters, and as a general rule an increase in absorption leads to a decrease in singularity density. Higher order singularities, which are the degeneracies of fundamental singularities, are not topologically stable and we do not see any universal statistical properties for them. Several questions about how the singularity abundance depends on absorption have been raised by these results and are suggested as future work. The statistical superuniversality seen here may depend on having perfect coupling as was found for complex time delay PDFs in Ref. [53]. How coupling between a scattering environment and the channels leading in and out affects the universal power laws, and similarly effects on the singularity density presented in this paper, are other unexplored problems. There are also opportunities for similar analysis in other fields of physics, as complex scalar fields or order parameters are commonly studied quantities.
Acknowledgements We acknowledge helpful discussions with Thomas Antonsen. We thank David Shrekenhamer and Timothy Sleasman of JHU APL for the design and fabrication of the metasurfaces used in this work. This work was supported by NSF/RINGS under grant No. ECCS-2148318, ONR under grant N000142312507, DARPA WARDEN under grant HR00112120021.
@@ Appendix A Simple Mathematical Reason for Universal Power Laws
[…]
@@
[…]
which has the empirically found $-3$ power law. Something similar can be done to show that the PDF $\mathcal{P}_{w}(w)$ has a $-2$ power law, where $w=\frac{1}{u}$. The steps are the same:
[…]
@@
[…]
and results in the expected $-2$ power law.
@@
Figure 7: PDFs of the inverse of complex Gaussian random numbers $\mathcal{G_{N}}$. These PDFs show the same power law tails as the experimental data and RMT numerics: $-3$ for $\mathcal{P}(\frac{\langle|\mathcal{G_{N}}|\rangle}{|\mathcal{G_{N}}|})$ and $-2$ for $\mathcal{P}(\frac{\langle|\mathrm{Re}[\mathcal{G_{N}}]|\rangle}{|\mathrm{Re}[\mathcal{G_{N}}]|})$ and $\mathcal{P}(\frac{\langle|\mathrm{Im}[\mathcal{G_{N}}]|\rangle}{|\mathrm{Im}[\mathcal{G_{N}}]|})$.
@@
To show this is purely a mathematical result which requires no physics, we generated $10^{9}$ complex numbers $\mathcal{G_{N}}$ from a Gaussian normal distribution. Fig. 7 shows the PDFs of $\frac{\langle|\mathcal{G_{N}}|\rangle}{|\mathcal{G_{N}}|}$, $\frac{\langle|\mathrm{Re}[\mathcal{G_{N}}]|\rangle}{|\mathrm{Re}[\mathcal{G_{N}}]|}$, and $\frac{\langle|\mathrm{Im}[\mathcal{G_{N}}]|\rangle}{|\mathrm{Im}[\mathcal{G_{N}}]|}$. Indeed, we see the same $-3$ and $-2$ power laws as in our experimental scattering matrix data.
Appendix B Singularities in $d=3$ Parameter Space

### v3

@@ III.1 Fundamental Singularities - Statistics
In panels (a-b) of Fig. 3, the color gradient corresponds to the absorption strength $\eta$, and there is a systematic, though small, vertical offset of the PDFs based on the ensemble value of $\eta$. This leads to different abundances of extreme values of $\mathcal{S}$. It is clear that the degree of loss is the strongest predictor of singularity density in parameter space. For any given quantity, the PDFs from ensembles with higher loss turn over to the power law tail at smaller values of the quantity and therefore have a lower overall probability at the extreme values associated with the scattering singularities. This agrees with the finding in [shaibe2025], that the time delay PDFs from ensembles with larger $\eta$ had the power law tails onset earlier, resulting in smaller probabilities at extreme values, hence fewer singularities in systems with more uniform absorption. Note, however, that this is only true for systems with sufficient absorption. We do not have experimental data from an ensemble with $\eta<1.8$, but since certain singularities are impossible in lossless systems (consider CPA which requires $\text{det}S=0+i0$, but if $\eta=0$ then $|\text{det}S|=1$) we expect there must be some range of $\eta$ where increasing loss increases the abundance of singularities.
@@ III.2 Singularity Dependence on Absorption
Figure 4: Singularity density as a function of uniform loss in (a) experimental ensembles and (b) RMT simulation. The overall trend of reduced singularity density with increasing loss in this finite range of $\eta$ is easier to see in (b) as the experimental systems have more differences than just $\eta$, which cannot all be normalized out.
It might seem counterintuitive that singularity abundance would decrease with increasing absorption since singularities are zeros of the complex wave fields and absorption should suppress the fluctuation magnitude of the waves. However, singularities, as stated before, are primarily the result of interference effects of superpositions of random waves. In systems with large dissipation, highly attenuated waves will not experience much destructive interference because the wave excitations become more homogeneous and less random.
In Fig. 4(a) we show the singularity mode density, defined as the number of singularities per mode in the ensemble, of the fundamental singularities EP-2, CPA, RSM-1, and TSM-1s (or equivalently TSM-2s for reciprocal $\beta=1$ systems) found in our experimental ensembles as a function of $\eta$. As discussed in Sec. II, the mean mode spacing $\Delta$ is system size and frequency dependent, so $\eta$ is correlated to $\mathcal{D}$ in our experimental systems. Also, the techniques used to create ensembles are different for each $\mathcal{D}$ so despite the great effort made to create diverse ensembles, there is some systematic dependence for how ‘ergodically’ we can explore the phase space of a given scattering environment based on what tunable parameters we have available. However, we still see a trend of lower singularity density with increasing $\eta$, agreeing with the complex time delay PDFs in Refs. [Lei2021Stats, shaibe2025] that downturn into the 3-3 tail at smaller time delay values for systems with larger $\eta$. We performed RMT simulations using a Gaussian Orthogonal Ensemble (GOE) of 600 random matrices of size $10^{5}\times 10^{5}$ to generate scattering matrices following the procedure detailed in Appendix E, Ref. [Lee2013] and Appendix A of Ref. [Hemmady2012]. We use the same set of Hamiltonians and vary only the loss $\eta$ between ensembles, resulting in Fig. 4(b).
@@
Figure 5: (a) Locations of scattering zeros in the ($\omega,\nu$) parameter space of an RMT simulation with $\eta=0$. (b) Locations of scattering zeros in the ($\omega,\nu$) parameter space of an RMT simulation using the same effective Hamiltonian and mode-channel coupling as in (a) but with $\eta=0.006$. Dashed red and black lines show the $\eta=0$ locations of TSM-2 and $\delta R_{12}=0$ which have become single (orange and black) points with the introduction of finite absorption. (c-d) Phase of $S_{+i}$ and $S_{-i}$ from simulation in (b). Black circles surround EPs (phase singularities) and arrows show direction of phase winding.
Note that at zero loss (i.e. a Hermitian system) certain singularities are forbidden, namely CPAs and EPs. Scattering parameter zeros (RSMs and TSMs) still occur because the elements of the $S$-matrix remain subunitary, as do symmetric reflection points and reciprocal points in non-reciprocal systems. However, the singularities can have different characteristics in a Hermitian system. Because we cannot experimentally measure a lossless system, we turn once more to RMT simulation. We consider the case of a reciprocal system with two channels ($\beta=1,M=2$) for simplicity using the model described in Ref. [erb2025] Appendix H. This model allows for Hamiltonian $H_{0}$ with a tunable parameter $\nu$ with the equation
@@
where $H_{1},H_{2}$, and $H_{3}$ are all standard GOE random matrices. A scattering matrix that depends on both the parameter $\nu$ and a frequency $\omega$ can be calculated using coupled mode theory [Fan2003]:
@@
We notice that in non-absorbing systems, what determines the abundance of singularities is more complicated than a single parameter ($\eta$) having the same effect on every singularity across different systems. Some singularities (e.g. CPAs and EPs) will not appear at all, while other singularities (e.g. symmetric reflection) may have certain properties (such as being a point or a curve in $d=2$ dimensional parameter space) depend on the number of channels $M$. The geometrical properties of the cavity, such as size and shape, can also determine if certain singularities are possible [Mostafazedah2011, Mostafazadeh2012, Joshi2024]. Because of the high degree of variability between the properties of different singularities at zero loss (or equivalently, balanced absorption and gain), no universal statements can be made about singularity statistics in unitary scattering systems.
In Fig. 5(b) we show the resulting singularities when we recalculate the $S$-matrix using the same model and parameters as in (a), but now with a small uniform attenuation $\eta=0.006$. We find that when a DP is perturbed in a way that introduces non-Hermiticity, it transforms into two EP-2s, similar to what was seen by Chen et al. in Ref. [Chen2025]. This is true for infinitesimally small amounts of loss, the smallest attempted in the simulations was $\eta=6\times 10^{-9}$, which can be seen in Appendix F. The charge of an EP-2 in a $M=2$ system is determined by whether the EP is a complex zero of $S_{+i}:=S_{11}-S_{22}-2i\sqrt{S_{12}S_{21}}$ (+i EPs marked in dark blue) or $S_{-i}:=S_{11}-S_{22}+2i\sqrt{S_{12}S_{21}}$ (-i EPs marked in light blue) [erb2025], the phase of which are shown in panels (c) and (d), respectively. The black circles and winding arrows show that each DP splits into a pair of EP-2s with opposite winding and charge $(\pm i$), such that the overall value of both these quantities is conserved when uniform loss is introduced.
In contrast, CPAs require some finite amount of loss to bring an $S$-matrix eigenvalue from the complex unit circle to zero. Exactly how much loss is required is a nuanced question since the trajectory of scattering eigenvalues in the complex plane is not linear. A lower bound, if it exists, should in principle be calculable, but is unknown to us. In the region of parameter space for the Hamiltonian used in Fig. 5, at $\eta=0.0006$ no CPAs could be found, however at $\eta=0.006$ a pair with opposite winding had already been created, marked by the green points in panel (b). With the addition of loss, not only have two CPAs appeared and the two DPs split into four EPs, but also the RSM-2s separated into RSM-1s (the pair corresponding to $|S_{22}|=0$ have already annihilated by $\eta=0.006$ so only two remain), and the $\delta R_{12}$ and TSM-2 lines have broken up into black and orange single points, respectively.
@@ III.3 Higher Order Singularities - Statistics
Finally, we consider the question of higher order singularities, which are composites of multiple fundamental singularities. Through the use of tunable perturbers, it is possible to make independent singularities coincident in parameter space, which results in novel applications such as the 50:50 IQ power splitter described in [erb2025] using a coincident CPA-EP-2. The intersection of complex zeros of different fields has no topological stability, and the singularities can move apart under arbitrarily small perturbation. Consequently, we do not find a universal power law for all PDFs of functions that diverges at a singularity of order $P>2$, for any given $P$.
Fig. 6(a) demonstrates the experimental PDFs of $\frac{\langle|\mathcal{S}|\rangle}{|\mathcal{S}|}$ for $\mathcal{S}=1-|C|+|\text{det}S|$ which diverge at a combined CPA-EP-2 of the $S$-matrix. The different curves come from 6 ensembles with different values for the parameters $\mathcal{D},M,\beta,\eta$. The bright pink curve, which comes from the lossiest system measured ($\eta=145$), has a power law fit of 3.49-3.49 as demonstrated by the dashed blue line, while the bright yellow curve which comes from the system with the least absorption ($\eta=1.8$) has a power law fit of 4.27-4.27, demonstrated by the dashed black line. The four other curves at intermediate values of uniform attenuation fill out the space in between with slopes bounded by those two values. This shows that even though $\mathcal{P}(\frac{\langle|\text{det}S|\rangle}{|\text{det}S|})$ and $\mathcal{P}(\frac{\langle 1-|C|\rangle}{1-|C|})$ have universal power laws, the power law of a PDF of a higher order combined singularity is system detail-dependent. It is interesting to note that the steepest power law comes from the ensemble with the lowest loss and increasing absorption makes the PDF drop slower. This implies that although the density of individual singularities decreases with increasing loss, their intersections become more common. It is currently unclear why this would be the case.
@@
A further statement should be made at this stage about higher order singularities of the scattering matrix. The fundamental singularities being combined are of not wholly independent complex scalars. CPAs are zeros of $\text{det}S$, but EPs also depend on the principal invariants of the $S$-matrix, one of which is its determinant. Similarly, $S_{xx}$ and $S_{xy}$ have inherent correlations which can depend on many factors, including the number of channels $M$ or loss $\eta$, as well as more complicated issues not considered in this paper such as imperfect mode-channel coupling. It is possible that truly universal statistics exist for composites of fully independent singularities, but it is very difficult to construct multiple complex scalar functions with no correlations out of a physical scattering matrix.
@@ IV Conclusion
We have examined the statistical properties of $S$-matrix singularities by generalizing the concept of singularity speckle patterns to arbitrary two dimensional parameter spaces and any complex scalar function that describes wave phenomena involving complicated scattering. The statistics discussed come from experimental microwave resonant systems with the number of channels ranging from $M=1$ to $M=4$. Both systems with and without Lorentz reciprocity were considered, as well as a wide range of uniform absorption ($1.8\leq\eta\leq 145$). Supporting RMT simulations were conducted with results that agree with the empirical data, showing these findings are generic to all wave scattering systems, such as acoustic, optical, and photonic resonators. Fundamental singularities, which are topologically-protected vortices of complex scalar fields in two-dimensional parameter space, have a “superuniversal” statistical rule that a function which has a simple pole at a singularity has a PDF with a -3 power law tail. These PDFs can then be used to estimate singularity abundance based on system parameters, and a general rule for absorbing systems is an increase in absorption leads to a decrease in singularity density. Higher order singularities, which are the degeneracies of fundamental singularities, are not topologically stable and we do not see any universal statistical properties for them.
Several questions about how the singularity abundance depends on absorption have been raised by these results and are suggested as future work. The statistical “superuniversality” seen here may depend on having perfect coupling as was found for complex time delay PDFs in Ref. [shaibe2025]. How coupling between a scattering environment and the channels leading in and out affects the universal power laws and singularity density presented in this paper, are other unexplored problems. There are also opportunities for similar analysis in other fields of physics, as complex scalar fields or order parameters are commonly studied quantities.
Acknowledgements We acknowledge helpful discussions with Thomas M. Antonsen. We thank David Shrekenhamer and Timothy Sleasman of JHU APL for the design and fabrication of the metasurfaces used in this work. This work was supported by NSF/RINGS under grant No. ECCS-2148318, ONR under grant N000142312507, and DARPA WARDEN under grant HR00112120021.
@@ Appendix A Simple Mathematical Reason for Universal Power Laws
$\mathcal{P}_{\rho}(\rho)=\int\mathcal{P}_{r}(r)*\delta\left(\rho-\frac{1}{r}\right)*rdr.$ (6)
@@
$\begin{split}\mathcal{P}_{\rho}(\rho)&=\int\mathcal{P}_{r}(r)*\delta\left(r-\frac{1}{\rho}\right)\frac{1}{\rho^{2}}*rdr\\ &=\frac{1}{\rho^{3}}\mathcal{P}_{r}(\frac{1}{\rho})\end{split}$ (8)
which has the empirically found 3-3 power law. Something similar can be done to show that the PDF $\mathcal{P}_{w}(w)$ has a 2-2 power law, where $w=\frac{1}{u}$. The steps are the same:
$\mathcal{P}_{w}(w)=\int\mathcal{P}_{u}(u)*\delta\left(w-\frac{1}{u}\right)*du$ (9)
@@
$\begin{split}\mathcal{P}_{w}(w)&=\int\mathcal{P}_{u}(u)*\delta\left(u-\frac{1}{w}\right)\frac{1}{w^{2}}*du\\ &=\frac{1}{w^{2}}\mathcal{P}_{u}(\frac{1}{w})\end{split}$ (11)
and results in the expected 2-2 power law.
@@
Figure 7: PDFs of the inverse of complex Gaussian random numbers $\mathcal{G_{N}}$. These PDFs show the same power law tails as the experimental data and RMT numerics: 3-3 for $\mathcal{P}(\frac{\langle|\mathcal{G_{N}}|\rangle}{|\mathcal{G_{N}}|})$ and 2-2 for $\mathcal{P}(\frac{\langle|\mathrm{Re}[\mathcal{G_{N}}]|\rangle}{|\mathrm{Re}[\mathcal{G_{N}}]|})$ and $\mathcal{P}(\frac{\langle|\mathrm{Im}[\mathcal{G_{N}}]|\rangle}{|\mathrm{Im}[\mathcal{G_{N}}]|})$.
@@
To show this is purely a mathematical result which requires no physics, we generated $10^{9}$ complex numbers $\mathcal{G_{N}}$ from a Gaussian normal distribution. Fig. 7 shows the PDFs of $\frac{\langle|\mathcal{G_{N}}|\rangle}{|\mathcal{G_{N}}|}$, $\frac{\langle|\mathrm{Re}[\mathcal{G_{N}}]|\rangle}{|\mathrm{Re}[\mathcal{G_{N}}]|}$, and $\frac{\langle|\mathrm{Im}[\mathcal{G_{N}}]|\rangle}{|\mathrm{Im}[\mathcal{G_{N}}]|}$. Indeed, we see the same 3-3 and 2-2 power laws as in our experimental scattering matrix data.
Appendix B Other Power Law Tails
As discussed in the paper, the condition for the PDF of a quantity to have a 3-3 (2-2) power law tail is that the quantity diverges as a simple pole at a singularity (one degree of freedom constraint of a complex scalar function). If the divergence does not have the form of a simple pole, the tail of the PDF will not have a 3-3 power law, in general.
Figure 8: PDFs of one over one minus the eigenvector coalescence $(1-|C|)^{-1}$, Petermann factor $K$, and inverse condition number $r^{-1}$ from the same experimental ensemble as panel (c) of Fig. 3 in the main text. All three of these quantities diverge solely at EP-2s. Dashed black lines characterize the power laws of the tail behaviors.
As mentioned before, the condition number $r$ has a branch point at an EP-2 since it can be written as $r=\sqrt{1-|C|^{2}}$ [Guo2023]. Therefore, even though $\frac{1}{r}$ diverges at an EP-2, the same as $\frac{1}{1-|C|}$ and the Petermann factor $K$, the PDF $\mathcal{P}(\frac{\langle r\rangle}{r})$ has a tail with a 5-5 power law, not 3-3. In Fig. 8, we show the experimental PDFs for these three quantities, where the data comes from the same $\mathcal{D}=2,M=2,\beta=1,\eta=40$ ensemble as panel (c) of Fig. 3. To further illustrate that the 5-5 power law results from the way $\frac{1}{r}$ diverges at an EP-2, we show the PDFs of $\frac{\langle|S_{11}|\rangle}{|S_{11}|}$ raised to different powers in Fig. 9. We label each PDF by the power law of the tail and find that $\mathcal{P}(\frac{\langle|S_{11}|^{1/2}\rangle}{|S_{11}|^{1/2}})$ has the same 5-5 power law as $\mathcal{P}(\frac{\langle r\rangle}{r})$. Notably, $\frac{1}{|S_{11}|^{1/2}}$ has a branch point at an RSM-1, just like $\frac{1}{r}$ at an EP-2. The red and purple curves ($\mathcal{P}(\frac{\langle|S_{11}|^{2}\rangle}{|S_{11}|^{2}})$ and $\mathcal{P}(\frac{\langle|S_{11}|^{3}\rangle}{|S_{11}|^{3}})$, respectively) further show that the PDFs of functions that have higher order poles at singularities also do not have 3-3 power law tails.
Figure 9: PDFs of $\frac{\langle|S_{11}|\rangle}{|S_{11}|}$ raised to different powers. The $S_{11}$ data comes from the same experimental ensemble as panel (c) of Fig. 3 in the main text. All the quantities diverge solely at $S_{11}=0+i0$ (RSM-1). Dashed black lines characterize the power laws of the tail behaviors.
Appendix C Singularities in $d=3$ Parameter Space